OpenStudy (lopezking1):

Two end points of a line segment are (2, 8) and (-4, -8). What are the coordinates of the point on the line through which its bisector passes? (5, -6) (3, 8) (-2, 0) (-1, 0)

OpenStudy (lopezking1):

I think its B

OpenStudy (lopezking1):

@melody16 you think you can help me lol :$

OpenStudy (mertsj):

The bisector passes through the midpoint. Here is the midpoint formula: \[(\frac{x _{1}+x _{2}}{2},\frac{y _{1}+y _{2}}{2})\]

OpenStudy (lopezking1):

\[\left( \frac{ 2+(-4) }{ 2 } \right)+\left(\begin{matrix}\frac{ 8+(-8) }{ 2 } \\ \end{matrix}\right)\]

OpenStudy (lopezking1):

, not +

OpenStudy (mertsj):

Simplify

OpenStudy (lopezking1):

\[\left(\begin{matrix}\frac{ -2 }{ 2} \\ \end{matrix}\right), 2\]

OpenStudy (mertsj):

\[(\frac{2+-4}{2},\frac{8+-8}{2})=(-1,0)\]

OpenStudy (lopezking1):

where does the -1 come from though

OpenStudy (mertsj):

What is 2-4?

OpenStudy (lopezking1):

-2

OpenStudy (mertsj):

What is -2/2 ?

OpenStudy (mertsj):